Hello,
I am trying to figure out how to find the solution for this equation.
2x-y+3z=1
-x+2y-z=-2
x+y+2z=-1
Can anybody help me solve this problem?
thanks

Answer by ankor@dixie-net.com(22740)

2x - y + 3z = 1
-x +2y - z =-2
x + y + 2z = -1
:
No
Multiply the last equation by -1 and add and see what happens
2x - y + 3z = 1
-x +2y - z = -2
-x - y - 2z = 1
--------------------adding eliminates all three variables